OpenStudy (anonymous):

Volume of H2O 200.0 mL H2O Mass of NaOH 2.535 g NaOH Initial temperature in calorimeter 24.2 °C Final temperature in calorimeter 27.8 °C Calculate the amount of energy involved in this dissolving process.

OpenStudy (zpupster):

q = m c (T2-T1) 200 mL H2O X 1 g/1 mL = 200 g H2O c = 4.184 J/gC T2 = 27.8 C T1 = 24.2 C Heat gained by H2O = 200 g X 4.18 J/gC X (27.8 - 24.2) = 3010 J